OpenStudy (anonymous):

Find the limit; lim x->2 from the left (3)/(x-2)

OpenStudy (anonymous):

nope, not sure how that will apply to this problem.

OpenStudy (anonymous):

I am looking at other similar example and I think the answer to this might be -infinity but how?

OpenStudy (anonymous):

\[\lim_{x\rightarrow 2^-}\frac{3}{x-2}\] If we approach from the left, this implies\[x<2\]Hence \[x-2<0\]Thus we can conclude \[\frac{1}{x-2}<0\] Therefore,\[\lim_{x\rightarrow 2^-}\frac{3}{x-2}\rightarrow -\infty\]

OpenStudy (dumbcow):

for these type of limits, the graph can also help x=2 is a vertical asymptote  From graph you can see that as you approach "x=2" from left the curve goes toward neg infinity